# Sublinear Data Structures for Nearest Neighbor in Ultra High Dimensions

### v1

@@ 5.2 Data Structure for $\ell_{2}$-metric with Near-linear Space and Query Time
6                   Add $b$ to $I$ with probability $p^{(b)}\triangleq\max_{(i,j)\in{n\choose 2}}\frac{|c_{i}^{(b)}-c_{j}^{(b)}|^{2}}{\|c_{i}-c_{j}\|_{2}^{2}}$;
@@ 18       Let $u\in\mathbb{R}^{I}$, where $\forall b\in I,u^{(b)}=q^{(b)}/\sqrt{p^{(b)}}$;
We show the proof in the following subsections in line with the proof structure described in Section 5. Although the proof at a high level follows the same structure as for the $\ell_{1}$ case, there are significant differences between the proofs themselves. The main differences are in the dimension reduction as well as in bounding the space and query complexities. Let $\delta^{\prime}=\delta/4$, where each good event described above happens with probability $1-\delta^{\prime}$.
@@ 5.2.1 Upper Bound for Estimate of Distance between Query and its Nearest Neighbor
$\mathbb{E}[\|r_{i}-u\|_{2}^{2}]=\mathbb{E}\left[\sum_{b\in I}\left(r_{i}^{(b)}-u^{(b)}\right)^{2}\right]=\mathbb{E}\left[\sum_{t\in[T],b\in[d]}\mathbbm{1}_{\{\text{in iteration $t$, we add $b$ to $I$\}}}\left(r_{i}^{(b)}-u^{(b)}\right)^{2}\right]$
$=\sum_{t\in[T],b\in[d]}p^{(b)}\left(r_{i}^{(b)}-u^{(b)}\right)^{2}=\sum_{t\in[T],b\in[d]}p^{(b)}\left(\frac{c_{i}^{(b)}}{\sqrt{p^{(b)}}}-\frac{q^{(b)}}{\sqrt{p^{(b)}}}\right)^{2}=\sum_{t\in[T],b\in[d]}\left(c_{i}^{(b)}-q^{(b)}\right)^{2}=T\|c_{i}-q\|_{2}^{2}.$
@@
$\|r_{i^{*}}-u\|_{2}^{2}\leq\frac{T}{\delta^{\prime}}\|c_{i^{*}}-q\|_{2}^{2}.\qed$
5.2.2 Coordinate Selection Preserves Distance Ratio
In this section, we show that, the following holds: for any query point $q\in\mathbb{R}^{n}$, whose nearest neighbor is $c_{i^{*}}$, with probability $1-2\delta^{\prime}$, for all $c_{i}\in C$ such that $\frac{\|c_{i}-q\|_{2}}{\|c_{i^{*}}-q\|_{2}}\geq 1+\epsilon$, the estimated ratio is significantly larger than one, that is $\frac{\|r_{i}-u\|_{2}}{\|r_{i^{*}}-u\|_{2}}\geq 1+\epsilon\delta^{\prime}/12$. This statement is formalized in Lemma 5.23.
@@ 5.2.2 Coordinate Selection Preserves Distance Ratio
Definition 5.19 (Good region and bad region).
Consider two points $c_{1},c_{2}\in\mathbb{R}^{d}$, and a query point $q\in\mathbb{R}^{d}$. We define the good region w.r.t. $(c_{1},c_{2},q)$ as
$\mathsf{Good}(c_{1},c_{2},q)=\{z\in[d]\mid\min\{c_{1}^{(z)},c_{2}^{(z)}\}-\frac{24}{\epsilon\delta^{\prime}}|c_{1}^{(z)}-c_{2}^{(z)}|\leq q^{(z)}\leq\max\{c_{1}^{(z)},c_{2}^{(z)}\}+\frac{24}{\epsilon\delta^{\prime}}|c_{1}^{(z)}-c_{2}^{(z)}|\}.$
@@
Definition 5.20 (Restricted distance).
@@
$\frac{1152}{\epsilon^{2}\delta^{\prime 2}}\left(\|b-a\|_{2}^{2}-\mathsf{dist}^{2}_{\mathsf{Good}(a,b,q)}(b,a)\right)=\sum_{z\in\mathsf{Bad}(a,b,q)}2\left(\frac{24}{\epsilon\delta^{\prime}}\right)^{2}|b^{(z)}-a^{(z)}|^{2}\leq\left(1+\frac{2}{\epsilon}\right)^{2}\|b-a\|_{2}^{2}.$
@@
$\frac{\mathsf{dist}^{2}_{\mathsf{Good}(a,b,q)}(b,a)}{\|b-a\|_{2}^{2}}\geq\frac{1143}{1152}\Rightarrow\frac{\mathsf{dist}_{\mathsf{Good}(a,b,q)}(b,a)}{\|b-a\|_{2}}\geq\frac{13}{16}.\qed$
@@
Notice that $\forall t\in[T],\mathbb{E}[\sum_{z\in\mathsf{Good}}P_{t,z}]=\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},q)$ and $\mathbb{E}[\sum_{z\in\mathsf{Good}}Q_{t,z}]=\mathsf{dist}^{2}_{\mathsf{Good}}(c_{j},q)$. Because
$\mathbb{E}[\sum_{z\in\mathsf{Good}}P_{t,z}]=\sum_{z\in\mathsf{Good}}\mathbb{E}[P_{t,z}]=\sum_{z\in\mathsf{Good}}p^{(z)}\cdot\left(\frac{q^{(z)}-c_{i}^{(z)}}{\sqrt{p^{(z)}}}\right)^{2}=\sum_{z\in\mathsf{Good}}|q^{(z)}-c_{i}^{(z)}|^{2}=\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},q).$
@@
Let $10<\epsilon_{2}=\epsilon/16<$ 1. Using Chernoff bound and Lemma 5.21, we have
$\Pr\left[\left|\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}-\mathbb{E}\left[\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}\right]\right|\geq\epsilon_{2}\mathbb{E}\left[\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}\right]\right]\leq 2\exp\left(-\frac{\epsilon_{2}^{2}T\mathbb{E}[\sum_{z\in\mathsf{Good}}P_{t,z}]}{3(1+\frac{24}{\epsilon\delta^{\prime}})^{2}\|c_{i}-c_{j}\|_{2}^{2}}\right)$
$=2\exp\left(-\frac{\epsilon_{2}^{2}T}{3(1+\frac{24}{\epsilon\delta^{\prime}})^{2}}\frac{\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},q)}{\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})}\frac{\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})}{\|c_{i}-c_{j}\|_{2}^{2}}\right)\leq 2\exp\left(-\frac{\epsilon_{2}^{2}T}{3(1+\frac{24}{\epsilon\delta^{\prime}})^{2}}\cdot\frac{1}{4}\cdot\frac{1143}{1152}\right)\leq\frac{\delta^{\prime}}{2n},$
@@
Case 1. If $\mathbb{E}[\sum_{z\in\mathsf{Good}}Q_{t,z}]\geq\frac{1}{16}\cdot\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})$, using Chernoff bound and Lemma 5.21, we have
@@
Case 2. If $\mathbb{E}[\sum_{z\in\mathsf{Good}}Q_{t,z}]<\frac{1}{16}\cdot\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})$. Then we have $\mathbb{E}[\sum_{z\in\mathsf{Good}}P_{t,z}]>(\frac{1}{2}-\frac{1}{16})\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})$, which gives that $\frac{\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},q)}{\mathsf{dist}^{2}_{\mathsf{Good}}(c_{j},q)}>7$. Using Chernoff bound and Lemma 5.21, we have
$\Pr\left[\left|\sum_{t\in[T],z\in\mathsf{Good}}Q_{t,z}-\mathbb{E}[\sum_{t\in[T],z\in\mathsf{Good}}Q_{t,z}]\right|>\frac{1}{16}T\cdot\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})\right]<2\exp\left(-\frac{T\cdot\frac{1}{16}\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})}{3(1+\frac{24}{\epsilon\delta^{\prime}})^{2}\|c_{i}-c_{j}\|_{2}^{2}}\right)$
$\leq 2\exp\left(-\frac{T}{48(1+\frac{24}{\epsilon\delta^{\prime}})^{2}}\frac{\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})}{\|c_{i}-c_{j}\|_{2}^{2}}\right)\leq\frac{\delta^{\prime}}{2n}.$
@@
The third item of the lemma is proved. ∎
@@
Proof.
We know that $\|r_{i^{*}}-u\|_{2}^{2}\leq\frac{T}{\delta^{\prime}}\|c_{i^{*}}-q\|_{2}^{2}$ holds with probability $1-\delta^{\prime}$ by 5.18. Thus, it is sufficient to show the statement of the lemma holds under this assumption with probability $1-\delta^{\prime}$.
@@
Applying 5.9, we prove that $\frac{g_{1}^{\prime}+b_{1}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}\geq 1+\frac{\epsilon\delta^{\prime}}{4}$, that is $\frac{\|r_{i}-u\|^{2}_{2}}{\|r_{i^{*}}-u\|^{2}_{2}}\geq 1+\frac{\epsilon\delta^{\prime}}{4}$. Thus we have $\frac{\|r_{i}-u\|_{2}}{\|r_{i^{*}}-u\|_{2}}\geq 1+\frac{\epsilon\delta^{\prime}}{12}$. ∎
5.2.3 $\ell_{2}$ Dimension Reduction Preserves the Distance Ratio
@@
Theorem 5.24 (distributional Johnson-Lindenstrauss Lemma).
@@
Proof.
Applying Theorem 5.24 with $m=O(\log(n/\delta^{\prime})/(\epsilon^{2}\delta^{\prime 2}))$, given any $v\in R$, it holds with probability $1-\delta^{\prime}/(n+1)$, that $\|Mv-Mu\|_{2}\in[1-\epsilon\delta^{\prime}/100,1+\epsilon\delta^{\prime}/100]\|v-u\|_{1}$. By union bound, with probability $1-\delta^{\prime}$, it holds for all $v\in R$. By Lemma 5.23 it holds with probability $1-2\delta^{\prime}$ that for all $i\in[n]$, if $\frac{\|c_{i}-q\|_{2}}{\|c_{i^{*}}-q\|_{2}}\geq 1+\epsilon$, then $\frac{\|r_{i}-u\|_{1}}{\|r_{i^{*}}-u\|_{1}}\geq 1+\epsilon\delta^{\prime}/12$, thus $\frac{\|Mr_{i}-Mu\|_{2}}{\|Mr_{i^{*}}-Mu\|_{2}}>1$ with probability $1-3\delta^{\prime}$, i.e. the algorithm will not output $i$ as the index of the approximate nearest neighbor. ∎
Remark 5.26.
Using approximate nearest neighbour technqiues, e.g. [IW18], can further reduce the space at the cost of slower query, i.e. reducing the power of $\frac{\log(n)}{\epsilon\delta}$ in space complexity.
5.2.4 Space and Query Complexity

### v2

@@ B.2 Data Structure for $\ell_{2}$-metric with Near-linear Space and Query Time
6                   Add $b$ to $I$ with probability $p^{(b)}\triangleq\max_{(i,j)\in{\binom{n}{2}}}\frac{|c_{i}^{(b)}-c_{j}^{(b)}|^{2}}{\|c_{i}-c_{j}\|_{2}^{2}}$;
@@ 18       Let $u\in\mathbb{R}^{I}$, where $\forall b\in I,u^{(b)}=q^{(b)}/\sqrt{p^{(b)}}$;
We show the proof in the following subsections in line with the proof structure described in Section 4. In particular, we show that each of the following good events, which we restate here for the sake of completeness, hold with probability at least $1-\delta^{\prime}$, where $\delta^{\prime}=\delta/4$.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
Although the proof at a high level follows the same structure as for the $\ell_{1}$ case, there are significant differences between the proofs themselves. The main differences are in the dimension reduction as well as in bounding the space and query complexities.
B.2.1 Upper Bound for Estimate of Distance between Query and its Nearest Neighbor
@@
$\mathbb{E}[\|r_{i}-u\|_{2}^{2}]=$ $\mathbb{E}\left[\sum_{b\in I}\left(r_{i}^{(b)}-u^{(b)}\right)^{2}\right]$
$=$ $\mathbb{E}\left[\sum_{t\in[T],b\in[d]}\mathbbm{1}_{\{\text{in iteration $t$, we add $b$ to $I$\}}}\left(r_{i}^{(b)}-u^{(b)}\right)^{2}\right]$
$=$ $\sum_{t\in[T],b\in[d]}p^{(b)}\left(r_{i}^{(b)}-u^{(b)}\right)^{2}=\sum_{t\in[T],b\in[d]}p^{(b)}\left(\frac{c_{i}^{(b)}}{\sqrt{p^{(b)}}}-\frac{q^{(b)}}{\sqrt{p^{(b)}}}\right)^{2}$
$=$ $\sum_{t\in[T],b\in[d]}\left(c_{i}^{(b)}-q^{(b)}\right)^{2}=T\|c_{i}-q\|_{2}^{2}.$
@@
$\|r_{i^{*}}-u\|_{2}^{2}\leq\frac{T}{\delta^{\prime}}\|c_{i^{*}}-q\|_{2}^{2}.$
B.2.2 Coordinate Selection Preserves Distance Ratio
In this section, we show that, the following holds: for any query point $q\in\mathbb{R}^{n}$, whose nearest neighbor is $c_{i^{*}}$, with probability $1-2\delta^{\prime}$, for all $c_{i}\in C$ such that $\frac{\|c_{i}-q\|_{2}}{\|c_{i^{*}}-q\|_{2}}\geq 1+\epsilon$, the estimated ratio is significantly larger than one, that is $\frac{\|r_{i}-u\|_{2}}{\|r_{i^{*}}-u\|_{2}}\geq 1+\epsilon\delta^{\prime}/12$. This statement is formalized in Lemma B.15.
@@ B.2.2 Coordinate Selection Preserves Distance Ratio
Definition B.9 (Good region and bad region).
Consider two points $c_{1},c_{2}\in\mathbb{R}^{d}$, and a query point $q\in\mathbb{R}^{d}$. We define the good region $\mathsf{Good}(c_{1},c_{2},q)$ w.r.t. $(c_{1},c_{2},q)$ as
$\{z\in[d]\mid\min\{c_{1}^{(z)},c_{2}^{(z)}\}-\frac{24}{\epsilon\delta^{\prime}}|c_{1}^{(z)}-c_{2}^{(z)}|\leq q^{(z)}\leq\max\{c_{1}^{(z)},c_{2}^{(z)}\}+\frac{24}{\epsilon\delta^{\prime}}|c_{1}^{(z)}-c_{2}^{(z)}|\}.$
@@
Definition B.10 (Restricted distance).
@@
$\frac{1152}{\epsilon^{2}\delta^{\prime 2}}\left(\|b-a\|_{2}^{2}-\mathsf{dist}^{2}_{\mathsf{Good}(a,b,q)}(b,a)\right)=\sum_{z\in\mathsf{Bad}(a,b,q)}2\left(\frac{24}{\epsilon\delta^{\prime}}\right)^{2}\left|b^{(z)}-a^{(z)}\right|^{2}$
$\leq\left(1+\frac{2}{\epsilon}\right)^{2}\|b-a\|_{2}^{2}.$
@@
$\frac{\mathsf{dist}^{2}_{\mathsf{Good}(a,b,q)}(b,a)}{\|b-a\|_{2}^{2}}\geq\frac{1143}{1152}\Rightarrow\frac{\mathsf{dist}_{\mathsf{Good}(a,b,q)}(b,a)}{\|b-a\|_{2}}\geq\frac{13}{16}.$
@@
Notice $\forall t\in[T],\mathbb{E}[\sum_{z\in\mathsf{Good}}P_{t,z}]=\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},q)$ and $\mathbb{E}[\sum_{z\in\mathsf{Good}}Q_{t,z}]=\mathsf{dist}^{2}_{\mathsf{Good}}(c_{j},q)$. Because
$\mathbb{E}[\sum_{z\in\mathsf{Good}}P_{t,z}]=$ $\sum_{z\in\mathsf{Good}}\mathbb{E}[P_{t,z}]$
$=$ $\sum_{z\in\mathsf{Good}}p^{(z)}\cdot\left(\frac{q^{(z)}-c_{i}^{(z)}}{\sqrt{p^{(z)}}}\right)^{2}=\sum_{z\in\mathsf{Good}}|q^{(z)}-c_{i}^{(z)}|^{2}=\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},q).$
@@
Let $10<\epsilon_{2}=\epsilon/16<$ 1. Using Chernoff bound and Lemma B.11, we have
$\Pr\left[\left|\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}-\mathbb{E}\left[\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}\right]\right|\geq\epsilon_{2}\mathbb{E}\left[\sum_{t\in[T],z\in\mathsf{Good}}P_{t,z}\right]\right]$
$\leq$ $2\exp\left(-\frac{\epsilon_{2}^{2}T\mathbb{E}[\sum_{z\in\mathsf{Good}}P_{t,z}]}{3(1+\frac{24}{\epsilon\delta^{\prime}})^{2}\|c_{i}-c_{j}\|_{2}^{2}}\right)$
$=$ $2\exp\left(-\frac{\epsilon_{2}^{2}T}{3(1+\frac{24}{\epsilon\delta^{\prime}})^{2}}\frac{\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},q)}{\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})}\frac{\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})}{\|c_{i}-c_{j}\|_{2}^{2}}\right)$
$\leq$ $2\exp\left(-\frac{\epsilon_{2}^{2}T}{3(1+\frac{24}{\epsilon\delta^{\prime}})^{2}}\cdot\frac{1}{4}\cdot\frac{1143}{1152}\right)\leq\frac{\delta^{\prime}}{2n},$
@@
Case 1. If $\mathbb{E}[\sum_{z\in\mathsf{Good}}Q_{t,z}]\geq\frac{1}{16}\cdot\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})$, using Chernoff bound and Lemma B.11, we have
@@
Case 2. If $\mathbb{E}[\sum_{z\in\mathsf{Good}}Q_{t,z}]<\frac{1}{16}\cdot\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})$. Then we have $\mathbb{E}[\sum_{z\in\mathsf{Good}}P_{t,z}]>(\frac{1}{2}-\frac{1}{16})\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})$, which gives that $\frac{\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},q)}{\mathsf{dist}^{2}_{\mathsf{Good}}(c_{j},q)}>7$. Using Chernoff bound and Lemma B.11, we have
$\Pr\left[\left|\sum_{t\in[T],z\in\mathsf{Good}}Q_{t,z}-\mathbb{E}[\sum_{t\in[T],z\in\mathsf{Good}}Q_{t,z}]\right|>\frac{1}{16}T\cdot\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})\right]$
$<$ $2\exp\left(-\frac{T\cdot\frac{1}{16}\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})}{3(1+\frac{24}{\epsilon\delta^{\prime}})^{2}\|c_{i}-c_{j}\|_{2}^{2}}\right)$
$\leq$ $2\exp\left(-\frac{T}{48(1+\frac{24}{\epsilon\delta^{\prime}})^{2}}\frac{\mathsf{dist}^{2}_{\mathsf{Good}}(c_{i},c_{j})}{\|c_{i}-c_{j}\|_{2}^{2}}\right)\leq\frac{\delta^{\prime}}{2n}.$
@@
The third item of the lemma is proved.
@@
Proof B.16.
We know that $\|r_{i^{*}}-u\|_{2}^{2}\leq\frac{T}{\delta^{\prime}}\|c_{i^{*}}-q\|_{2}^{2}$ holds with probability $1-\delta^{\prime}$ by 11. Thus, it is sufficient to show the statement of the lemma holds under this assumption with probability $1-\delta^{\prime}$.
@@
Applying Lemma B.13 on $c_{i}$ and $c_{i^{*}}$ gives that with probability $1-\delta^{\prime}/n$,
@@
Applying 9, we prove that $\frac{g_{1}^{\prime}+b_{1}^{\prime}}{g_{2}^{\prime}+b_{2}^{\prime}}\geq 1+\frac{\epsilon\delta^{\prime}}{4}$, that is $\frac{\|r_{i}-u\|^{2}_{2}}{\|r_{i^{*}}-u\|^{2}_{2}}\geq 1+\frac{\epsilon\delta^{\prime}}{4}$. Thus we have $\frac{\|r_{i}-u\|_{2}}{\|r_{i^{*}}-u\|_{2}}\geq 1+\frac{\epsilon\delta^{\prime}}{12}$.
B.2.3 $\ell_{2}$ Dimension Reduction Preserves the Distance Ratio
@@
Theorem B.17 (distributional Johnson-Lindenstrauss Lemma).
@@
Proof B.18.
Applying Theorem B.17 with $m=O(\log(n/\delta^{\prime})/(\epsilon^{2}\delta^{\prime 2}))$, given any $v\in R$, it holds with probability $1-\delta^{\prime}/(n+1)$, that $\|Mv-Mu\|_{2}\in[1-\epsilon\delta^{\prime}/100,1+\epsilon\delta^{\prime}/100]\|v-u\|_{1}$. By union bound, with probability $1-\delta^{\prime}$, it holds for all $v\in R$. By Lemma B.15 it holds with probability $1-2\delta^{\prime}$ that for all $i\in[n]$, if $\frac{\|c_{i}-q\|_{2}}{\|c_{i^{*}}-q\|_{2}}\geq 1+\epsilon$, then $\frac{\|r_{i}-u\|_{1}}{\|r_{i^{*}}-u\|_{1}}\geq 1+\epsilon\delta^{\prime}/12$, thus $\frac{\|Mr_{i}-Mu\|_{2}}{\|Mr_{i^{*}}-Mu\|_{2}}>1$ with probability $1-3\delta^{\prime}$, i.e. the algorithm will not output $i$ as the index of the approximate nearest neighbor.
Remark B.19.
Using approximate nearest neighbour technqiues, e.g. [25], can further reduce the space at the cost of slower query, i.e. reducing the power of $\frac{\log(n)}{\epsilon\delta}$ in space complexity.
B.2.4 Space and Query Complexity